OpenStudy (anonymous):

WILL GIVE MEDAL !!! What are the solutions of the inequality? |5x| *>=* 20 A) x *>=* -4 B) x *>=* 4 C) x *>=* 4 or x *<=* -4 D) x *>=* -4 and x *<=* 4

OpenStudy (studygurl14):

When you have an absolute value equation/inequality, you will need to split the equation into two.

OpenStudy (anonymous):

So it will or C or D ?

OpenStudy (studygurl14):

\(\large |5x|\ge 20\) becomes... \(\large 5x\ge 20\) or \(\large 5x\le 20\) Can you do the rest?

OpenStudy (studygurl14):

oops, sorry forgot a negative sign

OpenStudy (studygurl14):

\(\large 5x\ge 20\) or \(\large 5x\le -20\)

OpenStudy (anonymous):

yeah but thats not an answer choice ?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

r u following me @aleroth ?

OpenStudy (anonymous):

Uhh, not intentionally? :D Anyway, you need to simplify what StudyGurl14 wrote further. Divide 5 on both sides for both inequalities.

OpenStudy (studygurl14):

@KALIAN You have to solve for x. I didn't finish it.

OpenStudy (anonymous):

How do I do that ?

OpenStudy (studygurl14):

@aleroth said how.

OpenStudy (anonymous):